What is the maximum kinetic energy in eV of electrons ejected from sodium metal by 450-NM EM radiation, given that the binding energy is 2.28 eV

1
Expert's answer
2022-04-07T13:09:07-0400

The maximum kinetic energy of electrons ejected from metal

"K_{\max}=\frac{hc}{\lambda}-\varphi"

"K_{\max}=\frac{1239.8 }{\lambda\:(\rm nm)}-\varphi"

"K_{\max}=\frac{1239.8 }{450\:(\rm nm)}-2.28=0.475\:\rm eV"